# Danny Halawi4, Fred Zhang4, Philip E. Tetlock2,5

### v2

@@ 1 Introduction
Forecasting the future is a challenging but necessary task (Armstrong, 2001, Tetlock and Gardner, 2015). Economic forecasts influence investment and hiring decisions (Christensen et al., 2018). And forecasts of the Covid-19 pandemic in early 2020 prompted local lockdowns to slow the spread of the virus Adam (2020). However, human forecasting is often expensive, time-consuming, applicable only in specific domains, and subject to concerns about human biases. Motivated by these limitations, recent work explores the use of machine learning (ML) models, especially large language models (LLMs), in automated forecasting (Hendrycks et al., 2021, Zou et al., 2022, Yan et al., 2024, Pratt et al., 2024, Phan et al., 2024, Fluri et al., 2024, Halawi et al., 2024).
@@ 1.1 Related work
In addition, Schoenegger and Park (2023) and Abolghasemi et al. (2023) compare the accuracy of GPT-4 and other LLMs to human forecasters. Schoenegger et al. (2024b) found that an ensemble of $12$ LLMs rivaled human forecasts in a forecasting tournament in 2023 with a small number of questions, but the results are statistically imprecise. Unlike our work, that tournament was run only once and is no longer updated. Also, our much larger question set allows us to make precise statistical claims about the performance of LLMs relative to each other and relative to human performers.
@@ 2 Preliminaries
A forecasting question asks for a probabilistic prediction of a future event. A forecaster may assign probabilities to potential outcomes of the event. Forecasting platforms, including prediction markets, host a wide range of questions. Many questions are of public interest, such as “Who will be the next US president in the 2024 election?”
@@ 3 Benchmark, leaderboard, and datasets
3.1 Updating the question bank
In an automated process that runs nightly at 0:00 UTC, questions are added to the question bank, resolution values are updated, and metadata generated, as shown in Figure 1. Details follow.
Daily, 0:00 UTCUpdate QuestionsQuestion FilesQuestion FilesUpdate ResolutionsResolution FilesResolution FilesUpdate MetadataMetadata FileMetadata FileQuestionBank
Figure 1: The automated system that pulls questions from forecasting platforms and datasets for question generation from data providers.
3.1.1 Updating questions and resolution values
@@
3.1.2 Updating forecast question metadata
@@
3.2 Generating the question set
Every other Sunday at midnight UTC, we release a set of 1,000 forecast questions for LLMs through the process outlined in Figure 2. We sample an equal number of questions from each data source, to ensure representativeness. Within each source, we then uniformly sample questions across all question categories, aiming for an equal distribution from each category within each source. This ensures that models cannot be overfit to a specific type of question or topic. Limiting the number of questions generated to 1,000 also ensures that costs for testing LLM systems on the benchmark are capped for development teams with fewer resources.
For questions derived from dataset sources, the distribution of resolution dates is as follows, as represented by days past the forecast due date: [7, 30, 90, 180, 365, 1,095, 1,825, 3,650]. In other words, the distribution of resolution dates spans from one week into the future to ten years into the future. For questions derived from market sources, we ask only for the probability that each question will resolve positively (will the event underlying the question occur, or not occur). Prior to a given market question’s resolution, performance is evaluated by calculating the squared distance between the forecasted value and the crowd forecast on the platform from the previous day. In the main public leaderboard we report performance using all questions—including unresolved questions—but we also separately report performance on resolved questions for interested readers.
Every other Sunday,0:00 UTCSample QuestionsQuestionBankLLM Question Set   (1,000 Questions)Public Question Set(200 Questions)
Figure 2: We sample from the question bank every two weeks to produce question sets.
For every data source, there are two types of questions, each comprising half of the question set. The first type is a standard forecasting question with a binary outcome, e.g. “Will inflation (core CPI) be above 3% next month?” We construct the second type, combination questions, by pairing two standard questions. For combination questions, we ask for forecasts on all Boolean combinations of the two questions (i.e. $P(Q1\cap Q2)$, $P(Q1\cap\neg Q2)$, …). Considering the extensive number of existing standard questions and potential combinations in our question bank (as we could potentially combine 3, 4, or more standard questions together), we effectively have access to millions of possible forecasting questions to sample from. We show the number of two-question combination questions in the question bank as it stands at time of writing in the right-hand column of Table 1.
@@ 3.2 Generating the question set
To estimate baseline human and expert performance on these questions, we run surveys asking the general public and expert forecasters to predict on our question sets (see Section 4). The human question is comprised of 200 forecast questions sampled directly from the LLM set of 1,000 questions. [^5] We can evaluate humans and models on more than 200 questions because each data question asks about the value of a datapoint at several different time horizons; and we consider accuracy of the model or human forecaster at each time horizon to be a different forecast. We produce the question sets 10 days before the forecast due date to allow for time to create and run a human survey. LLMs receive their question set 24 hours before the due date, even though it was generated at the same time as the human question set. This gives us the ability to elicit forecasts, obtaining comparable forecast sets on the due date from both LLMs and humans who faced the same information environment. A detailed timeline is visible in Figure 3.
[…]
Figure 3: Timeline from question set generation to forecast due date, $D$.
@@ 3.3 Resolution
Markets
Ground truth is not available for market questions before the events are known to have occurred or not occurred. Before events resolve, we use the platform prediction (an aggregate of human forecasts reported on the platform) from the preceding day as a proxy resolution value for unresolved questions. These values are updated nightly as described in Section 3.1.1. This allows us to score the complete set of forecast questions, incorporating all information available to people on a nightly basis. Once a market question is officially resolved, we score the forecast against the resolution value, creating a definitive score for the question across all forecast horizons.
Datasets
Datasets can be updated as new information becomes available. Thus, questions derived from datasets are continuously resolved to the value from the the latest available data.
@@
For market questions, as we only ask users to forecast the outcome of the question, missing forecasts on question outcomes are assigned the value of the crowd forecast on the forecast due date, day $D$ in Figure 3. Some may argue this is overly-generous, but we did not want teams to have a competitive advantage by simply scraping the market websites themselves.
@@ 3.4 Leaderboards
We generate leaderboards, ranking models and humans by average overall score. The main leaderboard highlights the top forecasting submission across all questions and can by sorted by performance on the source type (market or dataset) and by score on resolved questions.
@@ 4.1 Public
We recruited 500 human forecasters via Prolific and Facebook advertising to participate as representatives of the general public. These human subjects completed a brief introductory survey to gather demographic information and evaluate performance on a few forecasting and comprehension tasks. We then drew 200 questions from the 1,000 question set provided to language models; human subjects completed a one-hour survey containing 20 random questions from this 200-item subset where they provided their forecasts and rationales for each question.
The number of responses per question in the 200-item subset varied to ensure representativeness across subject matters and sources; at least 40 responses were gathered per question, averaging 49 responses per question.
@@ 4.2 Superforecasters
We additionally recruited 39 expert forecasters with a strong track record of accurate performance on a diverse set of geopolitical questions (“superforecasters”) to participate in a 9-day forecasting experiment in which participants were prompted to give their individual forecasts for 20 random questions from the same 200-item subset of forecasting questions we provided to the general public. Roughly halfway through the 9-day experiment, participants were moved into a group forecasting stage in which we allowed these expert forecaster participants to see one another’s forecasts and rationales and to communicate with one another about each question. They were also given the opportunity to forecast on questions beyond the 20 questions assigned to them individually.
@@ Results
We evaluate human performance on the forecasting tasks by comparing predictions to the ground truth on resolved questions and comparing to community aggregates on unresolved questions. We find that the median public survey participant had an overall Brier score of 0.107 on the 200-item subset we drew from the LLM Question Set with 1,000 questions. The median superforecaster participant had an overall Brier score of 0.093 on this same subset of forecasting questions. See Section 5 and Table 2 for a comparison between human performance and LLM performance, as well as Appendix H for a top 50 leaderboard of human and LLM performance.
@@ 5.1 Methodology
We evaluate a suite of instruction-following chat models without any additional fine-tuning (see Section 2 for details on the models). In each baseline, we prompt the model to generate a probabilistic forecast that the question will resolve to “Yes."
@@
We implement six baselines: (1) zero-shot prompting, (2) prompting with scratchpad instructions, (3) zero-shot prompting with crowd forecasts, (4) scratchpad prompting with crowd forecasts, (5) prompting with scratchpad instructions and retrieved news articles, and (6) aggregating predictions from multiple LLMs. Each baseline is described in more detail below.
@@
    The question sets we provide to LLMs contain what we term freeze values for market questions, which are just the crowd forecast on the market the day the question set was created, as decsribed in Section 3.2. Our third baseline is the zero-shot prompt with freeze values.
@@
    Our fourth baseline is the scratchpad prompt with freeze values.
@@
    Since LLMs’ knowledge are not continuously updated, it is important to provide LLMs with up-to-date information relevant to the question (Zou et al., 2022). Our fifth baseline, scratchpad with information retrieval, uses the same scratchpad prompts that also contain retrieved news articles.
    The retrieval system is the same as described in Halawi et al. (2024): an LLM generates search queries for a news API, filters articles for relevancy, and summarizes the articles.
@@
    The aggregated prediction of forecasters (“crowd" prediction) is a strong benchmark. For instance, Metaculus (2023) shows that an ensemble of all forecasters consistently outperforms using just the top 5, 10, …, 30 best forecasters (based on past scores). In our final baseline, we aggregate the predictions generated by LLMs into an LLM “crowd" prediction.
    To produce the LLM crowd forecast, we query GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-Pro. We use three prompts crafted by superforecasters who were asked to write prompts that would help an LLM produce accurate forecasts (Figure 8, Figure 9, Figure 10). This results in $3\times 3=9$ forecasts per question. We then show 3 LLM crowd baselines using the median, geometric mean, and geometric mean of log odds (Satopää et al., 2014). See details at Appendix E.
@@ 5.2 Results
In Panel A of Table 2, we show that superforecasters achieve an overall mean Brier score of $0.093$, significantly outperforming both the general public (Brier $=0.107$, $p<0.01$) and the top LLM performer on the 200-item subset (Claude 3.5 Sonnet: Brier $=0.111$, $p\leq 0.01$). [^6] See statistical note in Appendix G. The top-performing language models all had access to the (human) crowd forecast on market questions. The top-performing model without access to the crowd forecast on market questions was significantly less accurate than models with access to the human forecast. The comparison between humans and LLMs relies on the 200 questions forecasted by humans, which is a random sub-sample of the 1,000 questions in the question set provided to LLMs. [^7] Accuracy measures are based on more than 200 forecasts because human and LLM forecasters submitted multiple forecasts on each dataset question, one for each time horizon. The results presented here include forecasts over the 7-day and 30-day time horizons.
Comparing LLMs
Panel B excludes humans and evaluates LLMs on the entire question set ($N=1,000$ questions). Here we see a similar ranking of models, with GPT-4-Turbo now slightly outperforming Claude-3.5 Sonnet. Both of those models outperform GPT-4o and Gemini-1.5 Pro. As in Panel A, most of the top-performing models use the scratchpad prompt (Figure 7) and use as inputs the human crowd forecasts for market questions. Access to recent topical news related to the questions did not improve performance.
Table 2: Leaderboard with Top 10 Models
Panel A: Includes human comparison group and evaluates the set of 200 questions received by humans
@@
Superforecaster median forecast ForecastBench 0.135 0.051 0.093 [0.073, 0.112] 0%
Public median forecast ForecastBench 0.167 0.048 0.107 [0.09, 0.125] 0.009 30%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.155 0.067 0.111 [0.09, 0.132] 0.01 33%
GPT-4 OpenAI Freeze values Zero shot 0.169 0.056 0.113 [0.093, 0.132] 0.003 37%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.155 0.074 0.114 [0.09, 0.139] 0.008 36%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.16 0.07 0.115 [0.095, 0.135] 0.004 31%
GPT-4o OpenAI Freeze values Scratchpad 0.173 0.059 0.116 [0.098, 0.133] 0.002 30%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.177 0.055 0.116 [0.097, 0.135] <0.001 37%
GPT-4o OpenAI News with freeze values Scratchpad 0.179 0.059 0.119 [0.101, 0.137] <0.001 42%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.163 0.081 0.122 [0.107, 0.137] <0.001 29%
Always 0.5 ForecastBench 0.25 0.165 0.208 [0.198, 0.218] <0.001 17%
Panel B: Excludes human comparison group and evaluates the entire set of 1,000 questions
@@
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.175 0.051 0.113 [0.107, 0.118] 0%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.179 0.047 0.113 [0.108, 0.119] 0.442 53%
GPT-4o OpenAI Freeze values Scratchpad 0.201 0.039 0.12 [0.115, 0.125] <0.001 45%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.168 0.075 0.121 [0.116, 0.127] <0.001 42%
GPT-4o OpenAI News with freeze values Scratchpad 0.196 0.05 0.123 [0.117, 0.128] <0.001 41%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.192 0.061 0.127 [0.12, 0.133] <0.001 44%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.172 0.082 0.127 [0.121, 0.132] <0.001 41%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.199 0.055 0.127 [0.12, 0.133] <0.001 47%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.176 0.078 0.127 [0.122, 0.133] <0.001 32%
GPT-4-Turbo-2024-04-09 OpenAI No Scratchpad 0.175 0.083 0.129 [0.124, 0.134] <0.001 7%
Always 0.5 ForecastBench 0.25 0.16 0.205 [0.202, 0.208] <0.001 22%
@@
1. We only show the performance of one model from each model series. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the static version presented here.
@@
As a particular failure mode, we find LLMs are specifically worse at combination questions. Although our human surveys did not explicitly ask for forecasts on combination questions, we bound human performance by assuming independence of the component of each combination question. This underestimates human accuracy because a human forecaster predicting the outcome of a combination question could account for independence between the permuted events. In Table 19, we present this comparison of human and LLM forecasts. We see that LLMs perform poorly on these combination questions, and including them in the benchmark widens the gap between human and LLM performance: superforecasters (Brier score of $0.068$) outperform the general public (Brier score of $0.083$) and the top LLM (Claude 3.5, Brier score of $0.119$) significantly. To benchmark the size of this gap in performance, the $0.051$ Brier score gap in performance between superforecasters and Claude 3.5 (the top-performing LLM) is significantly larger than the $0.031$ gap in performance between Claude 3.5 and Claude 2.1.

### v3

@@ 1 Introduction
Forecasting the future is a challenging but important task (Armstrong, 2001, Tetlock and Gardner, 2015). Economic forecasts influence investment and hiring decisions (Christensen et al., 2018). And forecasts of the Covid-19 pandemic in early 2020 prompted local lockdowns to slow the spread of the virus Adam (2020). However, human forecasting is often expensive, time-consuming, applicable only in specific domains, and subject to concerns about human biases. Motivated by these limitations, recent work explores the use of machine learning (ML) models, especially large language models (LLMs), in automated forecasting (Hendrycks et al., 2021, Zou et al., 2022, Yan et al., 2024, Pratt et al., 2024, Phan et al., 2024, Fluri et al., 2024, Halawi et al., 2024).
@@ 1.1 Related work
In addition, Schoenegger and Park (2023) and Abolghasemi et al. (2023) compare the accuracy of GPT-4 and other LLMs to human forecasters. Schoenegger et al. (2024b) found that an ensemble of $12$ LLMs rivaled human forecasts in a forecasting tournament in 2023 with a small number of questions, limiting the study’s statistical power. Unlike our work, that tournament was run only once and is no longer updated. Also, our much larger question set allows us to make precise statistical claims about the performance of LLMs relative to each other and relative to human performers.
@@ 2 Preliminaries
A forecasting question asks for a probabilistic prediction of a future event. A forecaster may assign probabilities to potential outcomes of the event. Forecasting platforms, including prediction markets, host a wide range of questions. Many questions are of public interest, such as “Will a Democrat win the 2028 US presidential election?”
@@ 3 Benchmark, leaderboard, and datasets
3.1 Question bank
In an automated process that runs nightly at 0:00 UTC, questions are added to the question bank, resolution values are updated, and metadata generated. Details follow.
3.1.1 Questions and resolution values
@@
3.1.2 Question metadata
@@
3.2 Question sets
LLM question set
We release a set of 1,000 forecast questions for LLMs every other Sunday at midnight UTC. We sample an equal number of questions from each data source to ensure representativeness. Within each source, we then uniformly sample questions across all question categories, aiming for an equal distribution from each category within each source. This ensures that models cannot be overfit to a specific type of question or topic. Limiting the number of questions generated to 1,000 also ensures that costs for testing LLM systems on the benchmark are capped for development teams with fewer resources.
Human question set
The human question is comprised of 200 forecast questions sampled directly from the LLM question set. When sampling from the LLM question set, we do our best to maintain proportionality across question sources and across categories within each question source; this ensures the question set addresses as many domains as possible.
Forecast horizons
For questions derived from dataset sources, the distribution of resolution dates is the $forecast\_due\_date+n$ days, where $n\in\{7,30,90,180,365,1095,1825,3650\}$. In other words, for each dataset question we ask for 8 forecasts, differing only in their resolution date, ranging from one week to 10 years. For questions derived from market sources, we ask for only 1 forecast: the probability that each question will resolve positively (will the event underlying the question occur, or not). In time, this setup will allow us to evaluate both human and LLM forecasting performance over the short, medium, and long term.
Combination questions
For every data source, there are two types of questions, each comprising half of the question set. The first type is a standard forecasting question with a binary outcome, e.g. “Will inflation (core CPI) be above 3% next month?” We construct the second type, combination questions, by pairing two standard questions. For combination questions, we ask for forecasts on all Boolean combinations of the two questions (i.e. $P(Q1\cap Q2)$, $P(Q1\cap\neg Q2)$, …). Considering the extensive number of existing standard questions and potential combinations in our question bank (as we could potentially combine 3, 4, or more standard questions together), we effectively have access to millions of possible forecasting questions to sample from. We show the number of two-question combination questions in the question bank as it stands at time of writing in the right-hand column of Table 1. This setup implies that for market combination questions, each forecaster will provide 4 forecasts, whereas for dataset combination questions, each forecaster will provide 32 forecasts (4 for each boolean combination of $Q1$ and $Q2$ at each of the 8 forecast horizons).
@@ 3.2 Question sets
Of the 1,000 questions in the LLM question set, 500 are combination questions. Each combination question is composed of two standard questions from the same set. This means LLMs will also provide forecasts for the individual components of these combination questions, since they’re drawn from the existing standard questions. Importantly, none of the 200 questions in the human question set are combination questions. [^5] We exclude combination questions because expert human forecasts are expensive to generate and by focusing on forecasts of standard questions we maximize the relevance of these expert human forecasts.
Timeline of forecasting round
To estimate baseline human and expert performance on these questions, we run surveys asking the general public and expert forecasters to predict on our question sets (see Section 4). We produce the question sets 10 days before the forecast due date to allow for time to create and run a human survey. LLM teams receive their question set 24 hours before the due date, even though it was generated at the same time as the human question set; this constrained time frame gives teams less time to be able to game the system, or be dishonest participants on the leaderbord. We thus elicit forecasts, obtaining comparable forecast sets on the due date from both LLMs and humans who faced the same information environment.
@@ 3.3 Resolution
Evaluating performance on market questions
Ground truth is not available for market questions before the events are known to have occurred or not occurred. Before questions resolve, we evaluate performance by calculating the squared distance between the forecast value and the platform prediction (an aggregate of human forecasts reported on the platform) from the preceding day. These values are updated nightly as described in Section 3.1.1. This allows us to score the complete set of forecast questions, incorporating all information available to people on a nightly basis. Once a market question is officially resolved, we score the forecast against the resolution value, creating a definitive score for the question.
Evaluating forecasts on unresolved market questions to the crowd forecast provides a good estimate of forecast performance, as crowd forecasts tend toward ground truth as the resolution date approaches. This allows us to estimate forecast performance on a larger set of market questions, as we would otherwise have to wait until each market resolved.
Evaluating performance on dataset questions
Datasets can be updated as new information becomes available. Thus, questions derived from datasets are continuously resolved to the value from the the latest available data. As the resolution dates for dataset questions mentioned in earlier come due (the 7 day horizon to the 10 year horizon), a new round of forecasts is evaluated. We thus are able to evaluate forecasting performance over different time horizons.
@@
For market questions, as we only ask users to forecast the outcome of the question, missing forecasts on question outcomes are assigned the value of the crowd forecast on the forecast due date. Some may argue this is overly-generous, but we did not want teams to have a competitive advantage by simply scraping the market websites themselves.
@@ 3.4 Leaderboards
We generate leaderboards, ranking models and humans by average overall score. The main leaderboard highlights the top forecasting submission across all questions and can by sorted by performance on the source type (market or dataset) and by resolution status (resolved or unresolved questions). The leaderboard is updated nightly, given the latest data, market resolution values, and crowd forecasts.
@@ 4 Human forecaster baseline
To compare LLM forecasting performance to human performance, we ran surveys of two different groups: the general public and expert forecasters (superforecasters).
@@ 4.1 Public
We recruited 500 human forecasters via Prolific and Facebook advertising to participate as representatives of the general public. These human subjects completed a brief introductory survey to gather demographic information [^6] See Appendix I for an overview of public participant demographics. and evaluate performance on a few forecasting and comprehension tasks. They then completed a one-hour survey containing 20 random questions from the 200-item human question set described in Section 3.2, providing their forecasts and rationales for each question.
The number of responses per question varied to ensure representativeness across subject matters and sources; at least 40 responses were gathered per question, averaging 49 responses per question.
@@ 4.2 Superforecasters
We additionally recruited 39 superforecasters, who have a strong track record of accurate performance on a diverse set of geopolitical questions, to participate in a 9-day forecasting experiment in which participants were prompted to give their individual forecasts for 20 random questions from the same 200-item human question set described above. Roughly halfway through the 9-day experiment, participants were moved into a group forecasting stage in which we allowed them to see one another’s forecasts and rationales and to communicate about each question. They were also given the opportunity to forecast on questions beyond the 20 questions assigned to them individually.
@@ Results
We evaluate human performance on the forecasting tasks by comparing predictions to the ground truth on resolved questions and comparing to community aggregates on unresolved questions. We find that the median public survey participant had an overall Brier score of 0.111 on the 200-item subset we drew from the LLM Question Set with 1,000 questions. The median superforecaster participant had an overall Brier score of 0.091 on this same subset of forecasting questions. See Section 5 and Table 2 for a comparison between human performance and LLM performance, as well as Appendix H for a top 50 leaderboard of human and LLM performance.
@@ 5.1 Methodology
We evaluate a suite of instruction-following chat models without any additional fine-tuning (see Section 2 for details on the models). In each baseline, we prompt the model to generate a probabilistic forecast that the question will resolve to “Yes.”
@@
We implement seven baselines: (1) zero-shot prompting; (2) prompting with scratchpad instructions; (3) prompting with scratchpad instructions and retrieved news articles; (4) zero-shot prompting with crowd forecasts; (5) scratchpad prompting with crowd forecasts; (6) scratchpad prompting with retrieved news articles and crowd forecasts; and (7) aggregating predictions from multiple LLMs. Each baseline is described in more detail below.
@@
    Since LLMs’ knowledge are not continuously updated, it is important to provide LLMs with up-to-date information relevant to the question (Zou et al., 2022). Our third baseline, scratchpad with news, uses the same scratchpad prompt as above, supplemented with retrieved news articles. The retrieval system is the same as described in Halawi et al. (2024): an LLM generates search queries for a news API, filters articles for relevancy, and summarizes the articles.
@@
    The question sets we provide to LLMs contain what we term freeze values. For market questions these are just the crowd forecast on the market the day the question set was created, as described in Section 3.2. For dataset questions, these are baseline values relevant to the forecasting task. [^7] For example, in a question generated from a Wikipedia page about whether a chess player’s Elo rating will increase by a given date, the freeze value is the chess player’s Elo rating on the question set generation date. An explanation of what the freeze value represents is also provided. Our third baseline is the zero-shot with freeze values. This is simply the zero-shot prompt from Baseline 1 supplemented with the freeze value and an explanation of the freeze value. For examples of the freeze value its explanation, see Table 8 and Table 9.
@@
    Our fifth baseline is the scratchpad with freeze values (the scratchpad prompt from Baseline 2 supplemented with freeze values as explained in Baseline 4).
@@
    Our sixth baseline is the scratchpad with news with freeze values.
  • 7
    The aggregated prediction of forecasters (“crowd” prediction) is a strong benchmark. For instance, Metaculus (2023) shows that an ensemble of all forecasters consistently outperforms using just the top 5, 10, …, 30 best forecasters (based on past scores). In our final baseline, we aggregate the predictions generated by LLMs into an LLM “crowd” prediction.
    To produce the LLM crowd forecast, we use 3 models (GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-Pro) and 3 prompts crafted by superforecasters (Figure 6, Figure 7, and Figure 8). This results in 9 forecasts per question. We generate 3 LLM crowd baselines using the median, geometric mean, and geometric mean of log odds (Satopää et al., 2014). For details, see Appendix E.
@@ 5.2 Results
In Table 2, we show that superforecasters achieve an overall mean Brier score of $0.092$, significantly outperforming both the general public (Brier $=0.114$, $p<0.001$) and the top LLM performer on the 200-item subset (Claude 3.5 Sonnet: Brier $=0.114$, $p=0.001$). [^8] See statistical note in Appendix G. The top-performing language models all had access to the crowd forecast on market questions (the “freeze values” from Baselines 4, 5, and 6 above). The top-performing model without access to the crowd forecast on market questions was less accurate than models with access to the human forecast with a Brier score of $0.126$. The comparison between humans and LLMs relies on the 200 questions forecasted by humans, which is a random sub-sample of the 1,000 questions in the question set provided to LLMs (excluding combination questions). [^9] Accuracy measures are based on more than 200 forecasts because human and LLM forecasters submitted multiple forecasts on each dataset question, one for each time horizon. The results presented here include forecasts over the 7-day, 30-day, and 90-day time horizons.
Table 2: LLM/Human Leaderboard (top 10)
@@
Superforecaster median forecast ForecastBench 0.126 0.059 0.092 [0.073, 0.112] 0%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.139 0.089 0.114 [0.092, 0.136] <0.01 31%
Public median forecast ForecastBench 0.156 0.072 0.114 [0.095, 0.133] <0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.144 0.094 0.119 [0.098, 0.141] <0.001 30%
GPT-4 OpenAI Freeze values Zero shot 0.164 0.079 0.121 [0.100, 0.143] <0.001 32%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.145 0.099 0.122 [0.095, 0.149] <0.001 32%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.170 0.077 0.124 [0.102, 0.145] <0.001 33%
GPT-4o OpenAI News with freeze values Scratchpad 0.168 0.082 0.125 [0.107, 0.144] <0.001 26%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.139 0.113 0.126 [0.104, 0.148] <0.001 29%
GPT-4o OpenAI Freeze values Scratchpad 0.163 0.093 0.128 [0.108, 0.148] <0.001 27%
Always 0.5 ForecastBench 0.250 0.184 0.217 [0.207, 0.227] <0.001 16%
Notes:
1. Shows performance on the 200 standard questions provided in the human question set at the 7-, 30-, and 90-day forecast horizons.
2. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the static version presented here.
3. For resolved questions, predictions are compared against ground truth while for unresolved questions, they are compared to community aggregates.
As a particular failure mode, we find LLMs are specifically worse at combination questions. Although our human surveys did not explicitly ask for forecasts on combination questions, we bound human performance by assuming independence of the component of each combination question. This underestimates human accuracy because a human forecaster predicting the outcome of a combination question could account for independence between the permuted events. In Table 20, we present this comparison of human and LLM forecasts. We see that LLMs perform poorly on these combination questions, and including them in the benchmark widens the gap between human and LLM performance: superforecasters (Brier score of $0.071$) outperform the general public (Brier score of $0.090$) and the top LLM (Claude-3.5 Sonnet, Brier score of $0.124$) significantly. To benchmark the size of this gap in performance, the $0.053$ Brier score gap in performance between superforecasters and Claude-3.5 Sonnet (the top-performing LLM) is significantly larger than the $0.035$ gap in performance between Claude-3.5 Sonnet and Claude 2.1.
Table 3: LLM Leaderboard (top 10)
@@
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.171 0.062 0.116 [0.111, 0.122] 0%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.171 0.065 0.118 [0.112, 0.123] 0.243 42%
GPT-4o OpenAI Freeze values Scratchpad 0.191 0.054 0.123 [0.117, 0.128] <0.001 43%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.162 0.089 0.125 [0.120, 0.131] <0.001 35%
GPT-4o OpenAI News with freeze values Scratchpad 0.192 0.066 0.129 [0.123, 0.134] <0.001 39%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.167 0.094 0.131 [0.125, 0.136] <0.001 34%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.190 0.076 0.133 [0.126, 0.139] <0.001 41%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.176 0.092 0.134 [0.128, 0.139] <0.001 31%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.171 0.097 0.134 [0.129, 0.139] <0.001 31%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.187 0.083 0.135 [0.129, 0.141] <0.001 32%
Always 0.5 ForecastBench 0.250 0.173 0.212 [0.209, 0.215] <0.001 25%
@@
1. Shows performance on the 1,000 (500 standard, 500 combination) questions in the LLM question set at the 7-, 30-, and 90-day forecast horizons.
2. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the static version presented here.
3. For resolved questions, predictions are compared against ground truth while for unresolved questions, they are compared to community aggregates.
@@
Comparing LLMs
Table 3 excludes humans and evaluates LLMs on the entire question set ($N=1,000$ questions). Here we see a similar ranking of models, with Claude-3.5 Sonnet slightly outperforming GPT-4-Turbo. Both of those models outperform GPT-4o and Gemini-1.5 Pro. As in Table 2, most of the top-performing models use the scratchpad prompt (Figure 5) and use as inputs the human crowd forecasts for market questions. Access to recent topical news related to the questions did not improve performance.
Refer to caption Chatbot Arena score (higher is better)Brier score (lower is better)
(a) Brier score vs. Chatbot Arena
Refer to caption Log training compute (FLOP)Brier score (lower is better)
(b) Brier score vs. Log training compute
Figure 1: Two graphs demonstrating the linear relationship between the Brier scores from Table 2 and (a) Chatbot Arena scores and (b) estimates of training compute. The dotted blue line represents the Superforecasters’ overall Brier score. A red dot with a bootstrapped 95% confidence interval is placed at the intersection of this dotted blue line with the dashed linear fit line to demonstrate the potential intersection of LLM Arena score/training compute and Superforecaster-level forecasting performance. For (b) , if estimates from Epoch AI (2024) were not available, we produced estimates following https://epoch.ai/blog/estimating-training-compute. The trend-line in (a) is $y=0.506-0.000298x$ ($R^{2}=0.47$) and in (b) it is $y=0.844-0.01213x$ ($R^{2}=0.41$).
LLM performance and forecasting accuracy
1(a) demonstrates the linear relationship between an LLM’s score on Chatbot Arena (see Chiang et al. (2024)) and the model’s overall Brier score from Table 2. We observe a significant correlation ($r=-0.69$, $p=0.002$), indicating that models with higher Arena scores also tend to produce more accurate forecasts. The calculated linear relationship implies that LLMs will match superforecaster performance when the Arena score approaches $1398$ (bootstrapped 95% CI: 13351335133513351581158115811581).
1(b) shows the log-linear relationship between the estimated training compute of LLMs and the model’s overall Brier score from Table 2. Projecting out the log-linear relationship, we find that LLMs could match superforecaster performance when training compute approaches $1\times 10^{27}$, though there is a large confidence interval (bootstrapped 95% CI: $7.08\times 10^{26}$$3.64\times 10^{30}$) because the log-linear relationship between training compute and forecasting accuracy is marginally statistically insignificant ($r=-0.64$, $p=0.06$).